OpenStudy (anonymous):

Given the equation (x + 3)2 + (y + 5)2 = 49 of a circle, identify the center and the radius

OpenStudy (anonymous):

do u know what's the standard eq for a circle?

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r

OpenStudy (anonymous):

it should be r^2...not just r. now, compare the 2 equations. what are the values of h,k and r?

OpenStudy (anonymous):

3 5 & 49

OpenStudy (anonymous):

don't ignore the signs. u should get, -h = 3 -k = 5 r^2 = 49. got it?

OpenStudy (anonymous):

ah ye ye i see

OpenStudy (anonymous):

right. so now can u tell me the values h, k and r? also, what do these variables represent in the standard eq?

OpenStudy (anonymous):

there are multiple choices right, two of them say 7 and the other say 49

OpenStudy (anonymous):

if r^2 = 49 then r = ?

OpenStudy (anonymous):

it says 7 or 49. its a ll confusing

OpenStudy (anonymous):

forget about those options. answer this: if u've been given x^2 = 100....then can u find the value of x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

seeing its a squared value, it'll be 10, no?

OpenStudy (anonymous):

that's right! now, a change in values.....given that r^2 = 49....what will be the value of r ?

OpenStudy (anonymous):

wouldnt it be 7..?

OpenStudy (anonymous):

bingo! so u know the values of h,k,r.....just list them for me. now, what do these variables represent, in the standard eq ?

OpenStudy (anonymous):

(x + 3)2 + (y + 5)2 = 49

OpenStudy (anonymous):

but its nit asking me for h,k,r.. just the C (circumfrence i guess) and r

OpenStudy (anonymous):

no. what i'm asking is: in the eq, (x-h)^2 + (y-k)^2 = r^2.........what do the variables h,k,r stand for?

OpenStudy (anonymous):

trust me. to find C(that's centre, not circumference) and r, u'll need to know h,k,r.

OpenStudy (anonymous):

mmmm thank you for helping but.. i figured it out. its the opposite of it, meaning the symbol of being positive or negative

OpenStudy (anonymous):

that's right. (h,k) = centre and r = radius.

OpenStudy (anonymous):

yea i see that.. thx

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

mm since yoyr like the first person who helps me, could you help me with the rest of my questions..?